Find the moles for 876 of He gas (22.4 liters/mole). Find the number of grams for 0.00176 moles CO2(6.022x10^23 molecules/mole)
Answer #1

no of moles of He   = W/G.A.Wt

                           = 876/4

                         = 219 moles

no of moles of CO2   = 0.00176moles

gram molar mass of CO2 = 44g/mole

mass of CO2   = no of moles * gram molar mass

                       = 0.00176*44

                        = 0.07744g >>>answer